James, a portfolio manager, would like to form the following portfolio between Microsoft and Coca-Cola:

            Expected return (%)      Standard Deviation (%) Weight

Microsoft                 28                          42                    0.4                  

Coca-Cola              12.5                         21                    0.6

The correlation between the two stocks is 0.5.

Now, suppose James can include the Treasury bills (T-bills) into his two-stock portfolio. The interest rate of the T-bills is 3.8%.

What is the Sharpe ratio of James’ portfolio?

For this question, please round up to the third decimal places.
Answer #1
Expected Return Standard Deviation Weight
Microsoft 28% 42% 0.4
Coca-Cola 12.50% 21% 0.6

Correlation between two stocks = ρ = 0.5

Weight of Microsoft in the portfolio = w1 = 0.4, Expected return on Microsoft = R1 = 28%, Standard Deviation of Microsoft = σ1 = 42%

Weight of Coca-Cola in the portfolio = w2 = 0.6, Expected return on Coca-Cola = R2 = 12.5%, Standard deviation of Coca-Cola = σ2 = 21%

Expected return of the portfolio is calculated using the formula: E[RP] = w1*R1 + w2*R2 = 0.4*28% + 0.6*12.5% = 18.7%

Variance of the portfolio is calculated using the formula:

σP2 = w1212 + w2222 + 2*ρ*w1*w212 = 0.42*(42%)2 + 0.62*(21%)2 + 2*0.5*0.4*0.6*(42%)*(21%) = 0.028224 + 0.015876 + 0.021168 = 0.065268

Standard deviation of the portfolio is square root of variance

Standard Devaition of the portfolio = σP = 0.0652681/2 = 0.255476026272525 = 25.5476026272525%

Risk-free rate = RF = 3.8%

Sharpe's ratio is calculated using the formula:

Sharpe's ratio = (E[RP] - RF)/σP = (18.7% - 3.8%)/25.5476026272525% = 0.583224978773768

Answer -> Sharpe's ratio of portfolio = 0.583